# Influence of Modeling Assumptions on the Inferred Dynamical State of Resonant Systems: A Case Study of the HD 45364 System

### v1

@@ 2.2 The influence of model priors and measurement noise on resonant dynamics
To further explore the influence of prior assumptions on the inferred dynamical state of the HD 45364 system, we run a second MCMC sampling simulation implementing the modified “resonance-weighted” prior, $\propto\exp\left[-\frac{1}{2S^{2}}A_{\mathrm{lib}}^{2}\right]$, with the value of $S$ set to $S=0.1$. The MCMC sampling was carried out with the same procedure described in Section 2.1, but using the modified log-likelihood function $\ln\mathcal{L}^{\prime}$ with $S=0.1$ instead of the unmodified log-likelihood function $\ln\mathcal{L}$. As in Section 2.1, $50$ chains were initialized in a multivariate Gaussian distribution centered on the best-fit parameters and each evolved for $50,000$ steps, generating a total of $2.5$ million posterior samples. We compute the resonant angles’ maximum deviation from equilibrium for $50,000$ random samples from the resonance-weighted prior model’s MCMC posterior using the same method as for the uninformative prior model, and plot their cumulative distributions as the dashed lines in Figure 3. We find that virtually all posterior samples generated with our modified priors have resonant angles that librate with small amplitude ($<90\degree$), demonstrating that our choice of model priors affects the inferred libration amplitudes.
@@
Figure 4: A contour plot of the log-likelihood $\ln\mathcal{L}$ of best-fit orbital solutions found by maximizing the modified log-likelihood, $\ln\mathcal{L}^{\prime}$, given by Equation (4), for different levels of the jitter scale factor, $f$, and libration amplitude penalty, $S$. Contours labeled $n\sigma$ mark the log-likelihood levels $\Delta\ln\mathcal{L}=-\frac{1}{2}n^{2}$, where $\Delta\ln{\mathcal{L}}$ is the log-likelihood function minus its global maximum value. Dashed contours have been linearly spaced at intervals of $0.25\sigma$ between $16\sigma$ and $17\sigma$ to show that low libration amplitude configurations may be erroneously ruled out compared to high libration amplitude configurations when instrumental jitter is not properly accounted for.
@@ 3.1 Resonance Capture
Tidal interactions between planets and a proto-planetary disk can cause their orbits to migrate and their eccentricities to damp (Lin & Papaloizou, 1979; Goldreich & Tremaine, 1980). Pairs of planets undergoing convergent migration will tend to capture into any first-order mean motion resonances they encounter, provided their migration rates are sufficiently slow (e.g., Murray & Dermott, 2000). After encountering a resonance, the planet pair will migrate together while maintaining a fixed period ratio and experience eccentricity growth until the resonant forcing of their eccentricities is counterbalanced by the damping effects of the disk (Snellgrove et al., 2001; Lee & Peale, 2002). For sufficiently smooth and slow migration forces, the pair of planets will eventually settle into an equilibrium configuration in which their resonant angles and eccentricities remain fixed. As discussed in Hadden & Payne (2020), for a given pair of planet masses and a specific $j$:$j-1$ resonance, these equilibrium configurations constitute a one-parameter family of orbital configurations that can be parameterized by the quantity
@@
where $e_{i}$ are the planets’ eccentricities, $\beta_{i}=m_{i}/(m_{1}+m_{2})$, $\alpha\approx\left(\frac{j-1}{j}\right)^{2/3}$ is the pair’s semi-major axis ratio, and $\Delta=\frac{j-1}{j}\frac{P_{2}}{P_{1}}-1$ measures the pair’s fractional deviation from exact period ratio commensurability.
@@
$\frac{d}{dt}\mathcal{D}\bigg{|}_{\mathrm{dis}}=0\\ \approx-\frac{\beta_{1}\sqrt{\alpha}e_{1}^{2}}{\tau_{e,1}}-\frac{\beta_{2}e_{2}^{2}}{\tau_{e,2}}-\frac{\beta_{1}\beta_{2}\sqrt{\alpha}}{3\left(j\beta_{1}\sqrt{\alpha}+(j-1)\beta_{2}\right)}\frac{3}{2\tau_{\alpha}}~{},$ (8)
@@
Figure 5 shows the eccentricities of a pair of planets in the family of 3:2 resonant equilibria. The planets’ masses are taken to be the median masses of HD 45364 b and c, as inferred from our MCMC simulations in Section 2.2. The equilibrium configurations are found by running a series of $N$-body simulations with eccentricity and semi-major axis damping forces included. We use the reboundx code’s (Tamayo et al., 2020) modify_orbits_direct effect to impose dissipative forces of the form of Equation (6) and (7). We set the planets’ eccentricity damping timescales to $\tau_{\mathrm{e,b}}=\tau_{\mathrm{e,c}}=:2\tau_{e}=10^{4}P_{b}$. The semi-major axis damping timescales, $\tau_{a,i}$, are chosen so that $K=\tau_{\alpha}/\tau_{e}$ ranges from 1 to $10^{3}$ and $\tau_{a,b}/\tau_{a,c}=-\frac{m_{b}}{m_{c}}\left(\frac{3}{2}\right)^{2/3}$. [^2] The latter condition is imposed for numerical convenience, as it ensures that the system remains at approximately the same mean stellocentric distance over the course of the simulations, obviating the need to adjust the simulation time step. The planets are initially placed in circular, coplanar orbits with a period ratio of $P_{c}/P_{b}=1.1\times\frac{3}{2}$, just outside the 3:2 MMR. The simulations are then integrated until the planets reach their equilibrium eccentricities. Plotting the equilibrium eccentricities of each simulation in Figure 5 produces a eccentricity “track,” along which each simulation evolves after capturing into resonance until it reaches its unique stopping point at which Equation (8) is satisfied (see Hadden & Payne, 2020).
@@ 3.2 Results
We determine possible migration scenarios for HD 45364 based on the posterior distribution of orbital configurations from the resonance-weighted prior model described in Section 2.2. Eccentricity posterior samples from our MCMC simulations with this model are plotted in Figure 5. Comparing the inferred distribution of the planets’ eccentricities to the range of equilibrium eccentricities generated by $N$-body simulations, we can estimate the relative strengths of eccentricity damping to orbital migration that led to the planets’ capture into the 3:2 MMR. The eccentricity posterior samples fall between the equilibria of simulations with $K\approx 12$ and $K\approx 175$, indicated by the red arrows in Figure 5.
@@
Figure 5: Comparison of our resonance capture simulations with the $2$-D marginalized distribution of planet eccentricities from the MCMC posterior of the resonance-weighted prior model described in Section 2.2 with $S=0.1$. The $1\sigma$, $2\sigma$ and $3\sigma$ contours of the posterior distribution are shaded. The simulated equilibrium eccentricity track using the best-fit planet masses is overlaid on the posterior distribution, with points falling within the MCMC posterior corresponding to simulated migration scenarios consistent with the observed eccentricities of the system. The approximate $K$ values for the bounds of the overlapping track segment up to the $3\sigma$ contours are labelled as well, suggesting the orbital configuration of HD 45634 b and c could be reproduced with a value $12\lesssim K\lesssim 175$.
@@
Finally, we also explicitly compute a corresponding $K$ value for each sample in the posterior distribution, assuming it is in an equilibrium MMR configuration. For each posterior sample, we substitute the planet masses and eccentricities into Equation (8) and solve for the value of $K=\tau_{\alpha}/\tau_{e}$ that gives $\frac{d}{dt}\mathcal{D}\bigg{|}_{\mathrm{dis}}=0$. While many posterior samples are not located exactly on the equilibrium line, the deviation is small so we nevertheless assign them the $K$ value of the associated MMR configuration. Computing the $K$ values of all posterior samples in this manner, we obtain the distribution shown in Figure 6. The $99\%$ credible interval of $12\lesssim K\lesssim 128$ for the $K$ distribution in Figure 6 is generally in agreement with the $3\sigma$ bounds estimated by comparing the posterior eccentricities to $N$-body simulations in Figure 5.
@@
Figure 6: Density histogram of $K$ values for the resonance-weighted prior model’s posterior eccentricity distribution shown in Figure 5, with the $68\%$, $95\%$ and $99\%$ credible intervals marked by red lines. The $K$ values are computed from the posterior distribution using the method described in Section 3.2. The histogram yields $12\lesssim K\lesssim 128$ as the $99\%$ credible interval for the $K$ values of possible migration scenarios, similar to the estimated bounds from the eccentricity track derived using $N$-body simulation in Figure 5.
4 Conclusions
@@
In this paper we re-analyzed the recently extended dataset of RV measurements from the planetary system HD 45364 and fit an RV signal to observations using an $N$-body dynamical model. We also investigated how the inferred dynamical state of the system can be influenced by the choice of model priors and by the assumed level of measurement noise. Finally, we explored possible scenarios for the migration history of HD 45364’s planets assuming a mean-motion resonance configuration of the system. Our conclusions are summarized as follows:
@@
    We demonstrate that model priors have a large influence on the inferred mean-motion resonance state of HD 45364 by comparing the results of an MCMC simulation that adopted a typical uninformative prior to one that adopted a prior more strongly weighted toward small libration amplitudes (i.e. resonant configurations). Furthermore, failure to properly account for instrumental jitter can lead to biased inference of libration amplitudes in general. In the case of HD 45364 specifically, orbital configurations with small libration amplitudes become disfavoured relative to those with large libration amplitudes if instrumental jitter is not appropriately accounted for.
@@
    We find that the current orbital configuration of the HD 45364 system is consistent with past smooth orbital migration and eccentricity damping resulting in capture into MMR. Furthermore, we estimate the ratio of migration to eccentricity damping timescales, $K=\tau_{a}/\tau_{e}$, during this process was in the range $10\lesssim K\lesssim 175$.
@@
    The large masses and closely spaced orbits of the planets in HD 45364 produce strong planet-planet gravitational interactions, which we leverage in our $N$-body dynamical model to break the $m\sin I$ degeneracy present in most RV observations. This allows us to constrain the masses of the planets to within a factor of $\sim 2$ of their median values and the inclination of the system to $\sin I\gtrsim 0.4$.
We envisage future application of the methods developed in this paper to other candidate systems with large planets in MMR (e.g. Wittenmyer et al., 2016; Rosenthal et al., 2019; Luque et al., 2019; Trifonov et al., 2019) in order to better understand the role of planetary migration in the formation history of these systems. In addition, we hope that our findings about the relationship between unmodelled RV measurement noise (instrumental jitter) and inferred libration amplitude will help inform studies of other resonant planetary systems.
@@ 5 Acknowledgments
We thank Hanno Rein for helpful discussions. S.H. acknowledges support by the Natural Sciences and Engineering Research Council of Canada (NSERC), funding references CITA 490888-16 and RGPIN-2020-03885.

### v2

@@ 2.2 The influence of model priors and measurement noise on resonant dynamics
To further explore the influence of prior assumptions on the inferred dynamical state of the HD 45364 system, we run a second MCMC sampling simulation implementing the modified “resonance-weighted” prior, $\propto\exp\left[-\frac{1}{2S^{2}}A_{\mathrm{lib}}^{2}\right]$, with the value of $S$ set to $S=0.1$. The MCMC sampling was carried out with the same procedure described in Section 2.1, but using the modified log-likelihood function $\ln\mathcal{L}^{\prime}$ with $S=0.1$ instead of the unmodified log-likelihood function $\ln\mathcal{L}$. As in Section 2.1, $50$ chains were initialized in a multivariate Gaussian distribution centered on the best-fit parameters and each evolved for $50,000$ steps, generating a total of $2.5$ million posterior samples. The chains successfully converged, with an estimated effective sample size of $\gtrsim 4000$. We compute the resonant angles’ maximum deviation from equilibrium for $50,000$ random samples from the resonance-weighted prior model’s MCMC posterior using the same method as for the uninformative prior model, and plot their cumulative distributions as the dashed lines in Figure 3. We find that virtually all posterior samples generated with our modified priors have resonant angles that librate with small amplitude ($<90\degree$), demonstrating that our choice of model priors affects the inferred libration amplitudes.
@@
Figure 4: A contour plot of the log-likelihood $\ln\mathcal{L}$ of best-fit orbital solutions found by maximizing the modified log-likelihood, $\ln\mathcal{L}^{\prime}$, given by Equation (4), for different levels of the jitter scale factor, $f$, and libration amplitude penalty, $S$. Contours labeled $n\sigma$ mark the log-likelihood levels $\Delta\ln\mathcal{L}=-\frac{1}{2}n^{2}$, where $\Delta\ln{\mathcal{L}}$ is the log-likelihood function minus its global maximum value. Dashed contours have been linearly spaced at intervals of $0.25\sigma$ between $13\sigma$ and $14\sigma$ to show that low libration amplitude configurations may be erroneously ruled out compared to high libration amplitude configurations when instrumental jitter is not properly accounted for.
@@ 3.1 Resonance Capture
Tidal interactions between planets and a proto-planetary disk can cause their orbits to migrate and their eccentricities to damp (Lin & Papaloizou, 1979; Goldreich & Tremaine, 1980). Pairs of planets undergoing convergent migration will tend to be captured into any first-order MMRs they encounter, provided their migration rates are sufficiently slow (e.g., Murray & Dermott, 2000). After encountering a resonance, the planet pair will migrate together while maintaining a fixed period ratio and experience eccentricity growth until the resonant forcing of their eccentricities is counterbalanced by the damping effects of the disk (Snellgrove et al., 2001; Lee & Peale, 2002). For sufficiently smooth and slow migration forces, the pair of planets will eventually settle into an equilibrium configuration in which their resonant angles and eccentricities remain fixed. As discussed in Hadden & Payne (2020), for a given pair of planet masses and a specific $j$:$j-1$ resonance, these equilibrium configurations constitute a one-parameter family of orbital configurations that can be parameterized by the quantity
@@
where $e_{i}$ are the planets’ eccentricities, $\beta_{i}=m_{i}/(m_{1}+m_{2})$, $\alpha\approx\left(\frac{j-1}{j}\right)^{2/3}$ is the pair’s semimajor axis ratio, and $\Delta=\frac{j-1}{j}\frac{P_{2}}{P_{1}}-1$ measures the pair’s fractional deviation from exact period ratio commensurability.
@@
$\frac{d}{dt}\mathcal{D}\bigg{|}_{\mathrm{dis}}$ $=0\approx-\frac{\beta_{1}\sqrt{\alpha}e_{1}^{2}}{\tau_{e,1}}-\frac{\beta_{2}e_{2}^{2}}{\tau_{e,2}}$
$-\frac{\beta_{1}\beta_{2}\sqrt{\alpha}}{3\left(j\beta_{1}\sqrt{\alpha}+(j-1)\beta_{2}\right)}\frac{3}{2\tau_{\alpha}}~{},$ (8)
@@
Figure 5 shows the eccentricities of a pair of planets in the family of 3:2 resonant equilibria. The planets’ masses are taken to be the median masses of HD 45364 b and c, as inferred from our MCMC simulations in Section 2.2. The equilibrium configurations are found by running a series of $N$-body simulations with eccentricity and semimajor axis damping forces included. We use the reboundx code’s (Tamayo et al., 2020) modify_orbits_direct effect to impose dissipative forces of the form of Equation (6) and (7). We set the planets’ eccentricity damping timescales to $\tau_{\mathrm{e,b}}=\tau_{\mathrm{e,c}}=:2\tau_{e}=10^{4}P_{b}$. The semimajor axis damping timescales, $\tau_{a,i}$, are chosen so that $K=\tau_{\alpha}/\tau_{e}$ ranges from 1 to $10^{3}$ and $\tau_{a,b}/\tau_{a,c}=-\frac{m_{b}}{m_{c}}\left(\frac{3}{2}\right)^{2/3}$. [^5] The latter condition is imposed for numerical convenience, as it ensures that the system remains at approximately the same mean stellocentric distance over the course of the simulations, obviating the need to adjust the simulation time step. The planets are initially placed in circular, coplanar orbits with a period ratio of $P_{c}/P_{b}=1.1\times\frac{3}{2}$, just outside the 3:2 MMR. The simulations are then integrated until the planets reach their equilibrium eccentricities. Plotting the equilibrium eccentricities of each simulation in Figure 5 produces an eccentricity “track,” along which each simulation evolves after capturing into resonance until it reaches its unique stopping point at which Equation (8) is satisfied (see Hadden & Payne, 2020).
@@ 3.2 Results
We determine possible migration scenarios for HD 45364 based on the posterior distribution of orbital configurations from the resonance-weighted prior model described in Section 2.2. Eccentricity posterior samples from our MCMC simulations with this model are plotted in Figure 5. Comparing the inferred distribution of the planets’ eccentricities to the range of equilibrium eccentricities generated by $N$-body simulations, we can estimate the relative strengths of eccentricity damping to orbital migration that led to the planets’ capture into the 3:2 MMR. The eccentricity posterior samples fall between the equilibria of simulations with $11\lesssim K\lesssim 144$, indicated by the red arrows in Figure 5.
@@
Figure 5: Comparison of our resonance capture simulations with the $2$D marginalized distribution of planet eccentricities from the MCMC posterior of the resonance-weighted prior model described in Section 2.2 with $S=0.1$. The $1\sigma$, $2\sigma$ and $3\sigma$ contours of the posterior distribution are shaded. The simulated equilibrium eccentricity track using the best-fit planet masses is overlaid on the posterior distribution, with points falling within the MCMC posterior corresponding to simulated migration scenarios consistent with the observed eccentricities of the system. The approximate $K$ values for the bounds of the overlapping track segment up to the $3\sigma$ contours are labeled as well, suggesting the orbital configuration of HD 45634 b and c could be reproduced with a value $11\lesssim K\lesssim 144$.
@@
Finally, we also explicitly compute a corresponding $K$ value for each sample in the posterior distribution, assuming it is in an equilibrium MMR configuration. For each posterior sample, we substitute the planet masses and eccentricities into Equation (8) and solve for the value of $K=\tau_{\alpha}/\tau_{e}$ that gives $\frac{d}{dt}\mathcal{D}\bigg{|}_{\mathrm{dis}}=0$. While many posterior samples are not located exactly on the equilibrium line, the deviation is small so we nevertheless assign them the $K$ value of the associated MMR configuration. Computing the $K$ values of all posterior samples in this manner, we obtain the distribution shown in Figure 6. The $99\%$ credible interval of $12\lesssim K\lesssim 117$ for the distribution of $K$ values in Figure 6 is generally in agreement with the $3\sigma$ bounds estimated by comparing the posterior eccentricities to $N$-body simulations in Figure 5.
@@
Figure 6: Density histogram of $K$ values for the resonance-weighted prior model’s posterior eccentricity distribution shown in Figure 5, with the $68\%$, $95\%$ and $99\%$ credible intervals marked by red lines. The $K$ values are computed from the posterior distribution using the method described in Section 3.2. The histogram yields $12\lesssim K\lesssim 117$ as the $99\%$ credible interval for the $K$ values of possible migration scenarios, similar to the estimated bounds from the eccentricity track derived using $N$-body simulation in Figure 5.
4 Discussion and Conclusions
@@
In this paper, we reanalyzed the recently extended data set of RV measurements from the planetary system HD 45364 and fit an RV signal to observations using an $N$-body dynamical model. We also investigated how the inferred dynamical state of the system can be influenced by the choice of model priors and by the assumed level of measurement noise. Finally, we explored possible scenarios for the migration history of HD 45364’s planets assuming an MMR configuration of the system.
Our conclusions are summarized as follows:
@@
    We demonstrate that model priors have a large influence on the inferred MMR state of HD 45364 by comparing the results of an MCMC simulation that adopted a typical uninformative prior to one that adopted a prior more strongly weighted toward small libration amplitudes (i.e., resonant configurations). Furthermore, failure to properly account for instrumental jitter can lead to biased inference of libration amplitudes in general. In the case of HD 45364 specifically, orbital configurations with small libration amplitudes become disfavoured relative to those with large libration amplitudes if instrumental jitter is not appropriately accounted for. More observations in the future will help conclusively determine whether the system is deep in the 3:2 MMR.
@@
    We find that the current orbital configuration of the HD 45364 system is consistent with past smooth orbital migration and eccentricity damping resulting in capture into MMR. Furthermore, we estimate the ratio of migration to eccentricity damping timescales, $K=\tau_{a}/\tau_{e}$, during this process was in the range $11\lesssim K\lesssim 144$.
@@
    The large masses and closely spaced orbits of the planets in HD 45364 produce strong planet-planet gravitational interactions, which we leverage in our $N$-body dynamical model to break the $m\sin I$ degeneracy present in most RV observations. This allows us to constrain the masses of the planets to within a factor of $\sim$1.51.51.51.5 of their median values and the inclination of the system to $\sin I\gtrsim 0.6$.
We emphasize that, at present, the data do not conclusively constrain the HD 45364 planets to be locked in the 3:2 MMR. Nonetheless, dynamical models with priors strongly weighted toward low libration amplitude configurations, the expected outcome of resonant capture via smooth migration, produce a quality of fit to the existing data that is nearly indistinguishable from the quality of fit obtained while adopting conventional uninformative priors. In principle, the full mathematical machinery of Bayesian model comparison could be deployed to ascribe a likelihood that the system’s planets reside in resonance. However, any quantitative determination of this likelihood would require apportioning prior probabilities to competing models with different assumptions about the system’s formation history. Since the current state of planet formation theory is far too uncertain to yield robust predictions for such prior probabilities, we have forgone any such quantitative Bayesian model comparison calculations. Instead, we simply note that the RV data for HD 45364 are wholly consistent with theoretical expectations for resonant planetary systems formed through a process of smooth, disk-induced migration. We envisage future application of the methods developed in this paper to other candidate systems with planets in or near MMR, including RV-detected systems (e.g. Wittenmyer et al., 2016; Luque et al., 2019; Rosenthal et al., 2019; Trifonov et al., 2019), transiting planet pairs (e.g. Dawson et al., 2021; Bozhilov et al., 2023) and higher-multiplicity resonant chains (e.g. Steffen et al., 2013; Mills et al., 2016; Luger et al., 2017; Leleu et al., 2021; MacDonald et al., 2021; Dai et al., 2023) in order to better understand the role of planetary migration in the formation history of these systems. In addition, we hope that our findings about the relationship between unmodeled RV measurement noise (instrumental jitter) and inferred libration amplitude will help inform studies of other resonant planetary systems.
@@ 5 Acknowledgments
We thank Hanno Rein for helpful discussions. We thank Trifon Trifonov for a careful referee report that helped improve this manuscript. S.H. acknowledges support by the Natural Sciences and Engineering Research Council of Canada (NSERC), funding references CITA 490888-16 and RGPIN-2020-03885.